# Verification and Synthesis of Compatible Control Lyapunov and Control Barrier Functions

### v1

@@ IV-A1 Exact Verification
$\Lambda(\mathbf{x})\mathbf{u}\leq\xi(\mathbf{x}),\ \forall\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$ (5)
@@
$\Lambda(\mathbf{x})=\begin{bmatrix}L_{g}V(\mathbf{x})\\ -L_{g}b(\mathbf{x})\\ A\end{bmatrix},\mathbf{\xi}(\mathbf{x})=\begin{bmatrix}-\kappa_{V}V(\mathbf{x})-L_{f}V(\mathbf{x})\\ \kappa_{b}b(\mathbf{x})+L_{f}b(\mathbf{x})\\ \mathbf{c}\end{bmatrix}.$
@@
The CBF $b(\mathbf{x})$ and CLF $V(\mathbf{x})$ are compatible if and only if the following set is empty:
$\{(\mathbf{x},\mathbf{z})\;|\;b(\mathbf{x})\geq 0,V(\mathbf{x})\leq 1,\\ \mathbf{z}^{T}\Lambda(\mathbf{x})=\mathbf{0}^{T},\mathbf{\xi}(\mathbf{x})^{T}\mathbf{z}=-1,\mathbf{z}\geq\mathbf{0}\},$ (6)
@@
The conclusion of Lemma 3 can be further simplified. Define the element-wise squared operation of vector $\mathbf{y}=[y_{1},\dots,y_{p+2}]^{T}$ as $\mathbf{y}^{2}=[y_{1}^{2},\dots,y_{p+2}^{2}]^{T}$. Then an equivalent formulation of Lemma 3 is as follows.
@@
The CBF $b(\mathbf{x})$ and CLF $V(\mathbf{x})$ are compatible if and only if the following set is empty:
$\{(\mathbf{x},\mathbf{y})\;|\;b(\mathbf{x})\geq 0,V(\mathbf{x})\leq 1,\\ (\mathbf{y}^{2})^{T}\Lambda(\mathbf{x})=\mathbf{0}^{T},\mathbf{\xi}(\mathbf{x})^{T}\mathbf{y}^{2}=-1\},$ (7)
@@
The given CLF $V(\mathbf{x})$ and CBF $b(\mathbf{x})$ are compatible if and only if there exist polynomials $\lambda(\mathbf{x},\mathbf{y}),\theta(\mathbf{x},\mathbf{y})$ satisfying:
@@
    $\theta(\mathbf{x},\mathbf{y})\in\Sigma[b(\mathbf{x}),1-V(\mathbf{x})]$
@@
s.t $\begin{split}\text{s.t }&-1-\mathbf{s}_{0}(\mathbf{x},\mathbf{y})^{T}((\mathbf{y}^{2})^{T}\Lambda(\mathbf{x}))-s_{1}(\mathbf{x},\mathbf{y})(\mathbf{\xi}(\mathbf{x})^{T}\mathbf{y}^{2}+1)\\ &\qquad-s_{2}(\mathbf{x},\mathbf{y})(1-V(\mathbf{x}))-s_{3}(\mathbf{x},\mathbf{y})b(\mathbf{x})\in\text{SOS}\end{split}$ (8b)
@@
A given pair of CLF $V(\mathbf{x})$ and CBF $b(\mathbf{x})$ are compatible if we can find the polynomials $s_{i}(\mathbf{x},\mathbf{y}),i=0,1,2,3$ in program (8).
@@ IV-A2 CBF Correctness Verification
In addition to verifying the CLF-CBF compatibility, we must also ensure that the set $\{\mathbf{x}\;|\;b(\mathbf{x})\geq 0\}$ is contained in the desired safe region, i.e., $\{\mathbf{x}|b(\mathbf{x})\geq 0\}$ implies $\mathbf{x}\in\mathcal{X}_{s}$ Assume an unsafe region presented by:
$\mathcal{X}_{u}=\cup_{j=1}^{K}\{\mathbf{x}|h_{j}(\mathbf{x})\leq 0\}.$ (9)
@@
s.t. $-1+q_{j}(\mathbf{x})b(\mathbf{x})+p_{j}(\mathbf{x})h_{j}(\mathbf{x})\in\text{SOS},\;j=1,\dots K.$ (10a)
@@
$\mathcal{X}_{u}=\cap_{j=1}^{K}\{\mathbf{x}|h_{j}(\mathbf{x})\leq 0\},$ (11)
@@
s.t. $-1+q(\mathbf{x})b(\mathbf{x})+\sum_{j=1}^{K}p_{j}(\mathbf{x})h_{j}(\mathbf{x})\in\text{SOS}.$ (12a)
If the unsafe region is only characterized by single polynomial, then (10) and (12) are the same and we would have single $q(\mathbf{x})$ and $p(\mathbf{x})$.
Both problems (10) and (12) can be solved
@@ IV-B Compatible CBF and CLF synthesis
Based on the previous verification framework, we next propose an alternating method to synthesize compatible CBF and CLF. We first formulate the synthesis problem, and then introduce the overall structure of the alternating algorithm and each component of the algorithm in detail. Finally we prove the compatibility of the synthesized CLF and CBF. For simplicity, in the following derivations, we use (9) as the definition of unsafe region and use the constraints in (10) as CBF correctness constraint.
@@
Given a system (1), with equilibrium state $\mathbf{x}^{*}=\mathbf{0}$, constants $\kappa_{V}$, $\kappa_{b}$, and polynomials $h_{1}(\mathbf{x})\dots h_{K}(\mathbf{x})$ characterizing the unsafe region, find a positive definite function $V(\mathbf{x})$ and a function $b(\mathbf{x})$ such that there always exist a $\mathbf{u}\in\mathcal{U}$ satisfying (4) for all the $\mathbf{x}\in\mathcal{D}\cap\mathcal{C}$, where $\mathcal{C}=\{\mathbf{x}\;|\;b(\mathbf{x})\geq 0\}$, and $\mathcal{D}=\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$. We call the region $\mathcal{C}\cap\mathcal{D}$ as the “compatible region”, and our goal is to maximize some objective function $\gamma(\mathcal{C}\cap\mathcal{D})$ on this compatible region. One example of the objective function is the volume $\gamma(\mathcal{C}\cap\mathcal{D})=\text{Vol}(\mathcal{C}\cap\mathcal{D})$.
@@
$\max_{\begin{subarray}{c}V(\mathbf{x}),b(\mathbf{x}),\\ p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})\end{subarray}}\gamma(\mathcal{C}\cap\mathcal{D})$ (13a)
@@
Unlike the verification problem (8) and (10) where $V(\mathbf{x})$ and $b(\mathbf{x})$ are given, now we need to jointly search for $V(\mathbf{x}),b(\mathbf{x})$ together with the Lagrangian multipliers $p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})$ in (13). Notice that all the constraints in problem (13) are convex, except for (8b) and (10a), which are bi-linear in the decision variable $V(\mathbf{x}),b(\mathbf{x}),p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})$, and hence cannot be directly handled by convex optimization. To remedy this, we resort to bilinear alternation, a common technique in SOS-based controller design [36, 41]. Where we alternate between
@@
    Lagrangian searching: fix $V(\mathbf{x}),b(\mathbf{x})$, search for the Lagrangian multipliers $p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})$.
@@
    Update CLF-CBF: fix the Lagrangian multiplier, search for $V(\mathbf{x}),b(\mathbf{x})$ while maximizing a surrogate function of $\gamma(\mathcal{C}\cap\mathcal{D})$.
Step 1 is the verification problem (8) and (10) discussed in the previous subsection. In step 2, we consider two different choices of objective function $\gamma(\mathcal{C}\cap\mathcal{D})$
  • Maximize the volume $\text{Vol}(\mathcal{C}\cap\mathcal{D})$.
  • Given a set of predefined “candidate states” $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(L)}$, we want the compatible region $\mathcal{C}\cap\mathcal{D}$ to cover as many “candidate states” as possible.
We next present how to search for $V(\mathbf{x}),b(\mathbf{x})$ to optimize these two objectives respectively, while fixing the Lagrangian multipliers.
IV-B1 Maximize volume
To maximize the volume of the compatible region, we need to solve the following optimization problem
$\max_{V(\mathbf{x}),b(\mathbf{x})}\;\text{Vol}(\mathcal{C}\cap\mathcal{D})$ (14a)
s.t $V(\mathbf{0})=0,V(\mathbf{x})\in\text{ SOS}$ (14b)
[…]
Directly maximing the volume of the semi-algebraic set $\text{Vol}(\mathcal{C}\cap\mathcal{D})$ can be challenging [^3] There exists some work to compute the volume of a given basic semi-algebraic set through a sequence of semidefinite programs with increasing sizes [42], but cannot be directly applied to maximize volume of the basic semi-algebraic set while searching for its defining polynomials., hence we seek an alternative approach to grow the compatible region, as shown in Appendix VII-A.
IV-B2 Cover many candidate states
Alternatively we can grow the compatible region $\mathcal{C}\cap\mathcal{D}$ to cover the candidate states $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(L)}$. We know that the candidate state $\mathbf{x}^{(i)}\in\mathcal{C}\cap\mathcal{D}$ iff $V(\mathbf{x}^{(i)})\leq 1$ and $b(\mathbf{x}^{(i)})\geq 0$. Hence we can design the objective function as
$\min_{V(\mathbf{x}),b(\mathbf{x})}\sum_{i=1}^{L}c_{1}\text{ReLU}(V(\mathbf{x}^{(i)})-1)+c_{2}\text{ReLU}(-b(\mathbf{x}^{(i))}),$ (15)
where $c_{1},c_{2}>0$ are given positive weights, and Eq (15) penalizes the violation of $x^{(i)}\notin\mathcal{C}\cap\mathcal{D}$. To avoid arbitrary scaling of CBF $b(\mathbf{x})$ (if $b(\mathbf{x})$ is a valid CBF, the scaling it as $k*b(\mathbf{x})$ by any positive constant $k$ is still a valid CBF, and the cost function becomes unbounded under infinite scaling), we impose an additional constraint $b(\mathbf{0})\leq 1$. As a result, the optimization problem is written as
@@ IV-B2 Cover many candidate states
$b(\mathbf{0})\leq 1$ (16d)
The optimization problems (14) and (16) are convex optimization problems when searching over $V(\mathbf{x}),b(\mathbf{x})$.
To summarize, we outline our synthesis algorithm in Algorithm 1
@@
1:Given system dynamics (1), positive constants $\kappa_{V},\kappa_{b}$ and unsafe region $\mathcal{X}_{u}$ in (9).
2:Initialize $V^{(0)}(\mathbf{x}),b^{(0)}(\mathbf{x})$. $i=0$
@@
4:     Fix $V^{(i)}(\mathbf{x}),b^{(i)}(\mathbf{x})$, solve (8) and (10) to find the Lagrangian multipliers $p^{(i)}(\mathbf{x}),q^{(i)}(\mathbf{x}),s^{(i)}(\mathbf{x},\mathbf{y})$.
5:     Fix $p^{(i)}(\mathbf{x}),q^{(i)}(\mathbf{x}),s^{(i)}(\mathbf{x},\mathbf{y})$, find CLF-CBF $V^{(i+1)}(\mathbf{x}),b^{(i+1)}(\mathbf{x})$ through solving either (14) or (16).
@@
In line 2 of Algorithm 1, we can initialize the CLF/CBF in many ways, for example a good candidate is to linearize the dynamics around the equilibrium state and use the LQR cost-to-go to design the initial guess.

### v2

@@ IV-A1 Compatibility Verification
$\Lambda(\mathbf{x})\mathbf{u}\leq\xi(\mathbf{x}),\ \forall\mathbf{x}\in\mathcal{C}\cap\mathcal{D},$ (5)
@@
$\Lambda(\mathbf{x})=\begin{bmatrix}L_{g}V(\mathbf{x})\\ -L_{g}h(\mathbf{x})\\ A\end{bmatrix},\mathbf{\xi}(\mathbf{x})=\begin{bmatrix}-\kappa_{V}(V(\mathbf{x}))-L_{f}V(\mathbf{x})\\ \kappa_{h}(h(\mathbf{x}))+L_{f}h(\mathbf{x})\\ \mathbf{c}\end{bmatrix}.$
@@
The CBF $h(\mathbf{x})$ and CLF $V(\mathbf{x})$ are compatible if and only if the following set is empty:
$\{(\mathbf{x},\mathbf{z})\;|\;h(\mathbf{x})\geq 0,V(\mathbf{x})\leq 1,\\ \mathbf{z}^{T}\Lambda(\mathbf{x})=\mathbf{0}^{T},\mathbf{\xi}(\mathbf{x})^{T}\mathbf{z}=-1,\mathbf{z}\geq\mathbf{0}\},$ (6)
@@
Lemma 3 can be further simplified. Define the element-wise squared operation of vector $\mathbf{y}=[y_{1},\dots,y_{p+2}]^{T}$ as $\mathbf{y}^{2}=[y_{1}^{2},\dots,y_{p+2}^{2}]^{T}$. Then an equivalent formulation of Lemma 3 is as follows:
@@
The CBF $h(\mathbf{x})$ and CLF $V(\mathbf{x})$ are compatible if and only if the following set is empty:
$\{(\mathbf{x},\mathbf{y})\;|\;h(\mathbf{x})\geq 0,V(\mathbf{x})\leq 1,\\ (\mathbf{y}^{2})^{T}\Lambda(\mathbf{x})=\mathbf{0}^{T},\mathbf{\xi}(\mathbf{x})^{T}\mathbf{y}^{2}=-1\},$ (7)
@@
The given CLF $V(\mathbf{x})$ and CBF $h(\mathbf{x})$ are compatible if and only if there exist polynomials $\lambda(\mathbf{x},\mathbf{y}),\theta(\mathbf{x},\mathbf{y})$ satisfying:
@@
    $\theta(\mathbf{x},\mathbf{y})\in\Sigma[h(\mathbf{x}),1-V(\mathbf{x})]$
@@
s.t $\begin{split}\text{s.t }&-1-\mathbf{s}_{0}(\mathbf{x},\mathbf{y})^{T}((\mathbf{y}^{2})^{T}\Lambda(\mathbf{x}))-s_{1}(\mathbf{x},\mathbf{y})(\mathbf{\xi}(\mathbf{x})^{T}\mathbf{y}^{2}+1)\\ &\qquad-s_{2}(\mathbf{x},\mathbf{y})(1-V(\mathbf{x}))-s_{3}(\mathbf{x},\mathbf{y})h(\mathbf{x})\in\text{SOS}\end{split}$ (8b)
@@
A given pair of CLF $V(\mathbf{x})$ and CBF $h(\mathbf{x})$ are compatible if we can find the polynomials $s_{i}(\mathbf{x},\mathbf{y}),i=0,1,2,3$ in program (8).
If the system dynamics also include the algebraic constraint $e(\mathbf{x})=0$ (namely (1c)), then we modify $\lambda(\mathbf{x},\mathbf{y})$ in Theorem 2 , by appending $e(\mathbf{x})$ to the ideal as $\lambda(\mathbf{x},\mathbf{y})\in\mathbb{I}[(\mathbf{y}^{2})^{T}\Lambda_{1}(\mathbf{x}),\dots,(\mathbf{y}^{2})^{T}\Lambda_{n_{u}}(\mathbf{x}),\mathbf{\xi}(\mathbf{x})^{T}\mathbf{y}^{2}+1,e(\mathbf{x})]$, and we change the S-procedure in (8) accordingly. For simplicity, we will ignore the system algebraic constraint $e(\mathbf{x})=0$ in the remaining of this methodology section, and we will provide concrete examples with such constraints in Section V for better understanding.
@@ IV-A2 CBF Correctness Verification
In addition to verifying the CLF-CBF compatibility, we must also ensure that the set $\{\mathbf{x}\;|\;h(\mathbf{x})\geq 0\}$ is contained in the desired safe region, i.e., $\{\mathbf{x}\;|\;h(\mathbf{x})\geq 0\}$ implies $\mathbf{x}\in\mathcal{X}_{s}$.
Assume an unsafe region presented as:
$\mathcal{X}_{u}=\cup_{j=1}^{K}\{\mathbf{x}|l_{j}(\mathbf{x})\leq 0\}.$ (9)
@@
s.t. $-1-q_{j}(\mathbf{x})h(\mathbf{x})+p_{j}(\mathbf{x})l_{j}(\mathbf{x})\in\text{SOS},\;j=1,\dots K.$ (10)
@@
$\mathcal{X}_{u}=\cap_{j=1}^{K}\{\mathbf{x}|l_{j}(\mathbf{x})\leq 0\},$ (11)
@@
s.t. $-1-q(\mathbf{x})h(\mathbf{x})+\sum_{j=1}^{K}p_{j}(\mathbf{x})l_{j}(\mathbf{x})\in\text{SOS}.$ (12)
Both problems (10) and (12) can be solved via SOS optimization.
@@ IV-B Compatible CBF and CLF synthesis
Based on the previous verification framework, we next propose an alternating method to synthesize compatible CBF and CLF. For simplicity, in the following derivations, we use (9) as the definition of unsafe region and use the constraints in (10) as CBF correctness constraint.
As mentioned previously in the end of Section III, we set the kappa function $\kappa_{V}(\cdot)$ and $\kappa_{h}(\cdot)$ as linear functions with positive scalar constant coefficient, and we synthesize CLF/CBF with exponential stability/safety. This choice of linear $\kappa$ functions renders (8b) to be a linear function of the CLF/CBF, allowing us to search for CLF/CBF through convex optimization.
@@
Given a system (1), with equilibrium state $\mathbf{x}^{*}=\mathbf{0}$, constants $\kappa_{V}$, $\kappa_{h}$, and polynomials $l_{1}(\mathbf{x})\dots l_{K}(\mathbf{x})$ characterizing the unsafe region, find a positive definite function $V(\mathbf{x})$ and a function $h(\mathbf{x})$ such that there always exist a $\mathbf{u}\in\mathcal{U}$ satisfying (4) for all the $\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$, where $\mathcal{C}=\{\mathbf{x}\;|\;h(\mathbf{x})\geq 0\}$, and $\mathcal{D}=\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$. We call the region $\mathcal{C}\cap\mathcal{D}$ as the “compatible region”, and our goal is to maximize some objective function $\gamma(\mathcal{C}\cap\mathcal{D})$ to expand this compatible region.
@@
$\max_{\begin{subarray}{c}V(\mathbf{x}),h(\mathbf{x}),\\ p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})\end{subarray}}\gamma(\mathcal{C}\cap\mathcal{D})$ (13a)
@@
Unlike the verification problem (8) and (10) where $V(\mathbf{x})$ and $h(\mathbf{x})$ are given, now we need to jointly search for $V(\mathbf{x}),h(\mathbf{x})$ together with the Lagrangian multipliers $p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})$ in (13). Notice that all the constraints in problem (13) are convex, except for (8b) and (10), which are bi-linear in the decision variable $V(\mathbf{x}),h(\mathbf{x}),p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x})$, and hence cannot be directly handled by convex optimization. To remedy this, we resort to bilinear alternation, a common technique in SOS-based controller design [37, 45]. Where we alternate between
@@
    Lagrangian searching: fix $V(\mathbf{x}),h(\mathbf{x})$, search for the Lagrangian multipliers $p(\mathbf{x}),q(\mathbf{x}),s(\mathbf{x},\mathbf{y})$.
@@
    Update CLF-CBF: fix the Lagrangian multiplier $q(\mathbf{x}),s(\mathbf{x},\mathbf{y})$, search for $V(\mathbf{x}),h(\mathbf{x})$ and $p(\mathbf{x})$ while maximizing a surrogate function $\gamma(\mathcal{C}\cap\mathcal{D})$.
Step 1 is the verification problem (8) and (10) discussed in the previous subsection. In step 2, we consider the following objective to expand the compatible region $\mathcal{C}\cap\mathcal{D}$: given a set of predefined “candidate states” $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(L)}$, we want the compatible region $\mathcal{C}\cap\mathcal{D}$ to cover as many “candidate states” as possible. We know that the candidate state $\mathbf{x}^{(i)}\in\mathcal{C}\cap\mathcal{D}$ iff $V(\mathbf{x}^{(i)})\leq 1$ and $h(\mathbf{x}^{(i)})\geq 0$. Hence we can design the objective function as minimizing
$\sum_{i=1}^{L}c_{1}\text{ReLU}\left(V(\mathbf{x}^{(i)})-1\right)+c_{2}\text{ReLU}\left(-h(\mathbf{x}^{(i)})\right),$ (14)
where $c_{1},c_{2}>0$ are given positive weights, and Eq (14) penalizes the violation $x^{(i)}\notin\mathcal{C}\cap\mathcal{D}$. To avoid arbitrary scaling of CBF $h(\mathbf{x})$ (if $h(\mathbf{x})$ is a valid CBF, the scaling it as $k*h(\mathbf{x})$ by any positive constant $k$ is still a valid CBF, and the cost function becomes unbounded under infinite scaling), we impose an additional constraint $h_{\text{lower}}\leq h(\mathbf{0})\leq h_{\text{upper}}$, where $h_{\text{lower}},h_{\text{upper}}$ are given positive constants. As a result, the optimization problem is written as
@@
$h_{\text{lower}}\leq h(\mathbf{0})\leq h_{\text{upper}}.$ (15d)
The optimization problem (15) is a convex optimization problem when searching over $V(\mathbf{x}),h(\mathbf{x})$ and $p(\mathbf{x})$.
To summarize, we outline our synthesis algorithm in Algorithm 1.
@@
1:Given system dynamics (1), positive constants $\kappa_{V},\kappa_{h}$ and unsafe region $\mathcal{X}_{u}$ in (9).
2:Initialize $V^{(0)}(\mathbf{x}),h^{(0)}(\mathbf{x})$. $i=0$
@@
4:     Fix $V^{(i)}(\mathbf{x}),h^{(i)}(\mathbf{x})$, solve (8) and (10) to find the Lagrangian multipliers $p^{(i)}(\mathbf{x}),q^{(i)}(\mathbf{x}),s^{(i)}(\mathbf{x},\mathbf{y})$.
5:     Fix $q^{(i)}(\mathbf{x}),s^{(i)}(\mathbf{x},\mathbf{y})$, find CLF-CBF $V^{(i+1)}(\mathbf{x}),h^{(i+1)}(\mathbf{x})$ and Lagrangian $p^{(i+1)}(\mathbf{x})$ through solving (15).
@@
In line 2 of Algorithm 1, we can initialize the CLF/CBF in many ways, for example a good candidate is to linearize the dynamics around the equilibrium state and use the LQR solution to design the initial guess (the LQR solution with the equality constraint $e(\mathbf{x})=0$ can be computed by projecting the dynamics onto the manifold $e(\mathbf{x})=0$, see [46, chapter 8.3.2]).